OpenStudy (anonymous):

What is the value of the expression? 6 • [(32 − 8) ÷ 4 + 2]

OpenStudy (anonymous):

@butterflydreamer

OpenStudy (butterflydreamer):

solve what's in the brackets first! (32 - 8) = ?

OpenStudy (onepieceftw):

Order of operations say to do parentheses first.

OpenStudy (anonymous):

24

OpenStudy (butterflydreamer):

Good and now 24 divided by 4 = ?

OpenStudy (anonymous):

= 6

OpenStudy (butterflydreamer):

correct! Now add 2 . 6 + 2 =?

OpenStudy (anonymous):

=8

OpenStudy (butterflydreamer):

yep and now, multiply by 6 and you're done :)

OpenStudy (anonymous):

can you help me with a few mor like this

OpenStudy (anonymous):

so the answer is 78

OpenStudy (butterflydreamer):

Sure. I've noticed that you've been posting similar questions though. Just remember the order of operations and BODMAS and you'll be fine :) And that is not quite the answer. 6 times 8 = ?

OpenStudy (anonymous):

oh = 48

OpenStudy (anonymous):

im sorry

OpenStudy (butterflydreamer):

that's correct :) Also, if you don't understand BODMAS, it's the same as PEMDAS http://prntscr.com/86271w <- super helpful reference with order of operations!

OpenStudy (anonymous):

ok so you can help me with two more qestions right

OpenStudy (onepieceftw):

BODMAS was before we added exponents to the mnemonic

OpenStudy (anonymous):

Which placement of grouping symbols in the expression gives it a value of 22? 3 • 4 – 2 + 5 +7 A. 3 • [4 – (2 + 5) + 7] B. (3 • 4) – [2 + (5 + 7)] C. [(3 • 4) – 2] + (5 + 7) D. 3 • (4 – 2) + (5 + 7)

OpenStudy (anonymous):

thats the nest one

OpenStudy (mathstudent55):

Try each choice using the correct order of operations, and see which one gives you the answer you need.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

What does the O mean in BODMAS?

OpenStudy (mathstudent55):

The way to do a problem like this is to do one step at a time and writing everything else each step. Of course, you need to follow PEMDAS for each step. 6 • [(32 − 8) ÷ 4 + 2] = = 6 • [24 ÷ 4 + 2] = 6 • [6 + 2] = 6 • 8 = 48 By writing everything each step, you don't forget any operation.

OpenStudy (mathstudent55):

Also, it is important to understand PEMDAS properly. Most people don't know how it works.

OpenStudy (mathstudent55):

@jpharper Let's work through each choice of the second problem.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

A. 3 • [4 – (2 + 5) + 7] = Innermost parentheses first: = 3 • [4 – 7 + 7] We still have the square parentheses, so we do that next.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

Inside the square brackets, we have a subtraction and an addition. In PEMDAS, AS means addition, Subtraction. Those two operations are the same level of priority, so we do them in the order they appear from left to right. = 3 • [-3 + 7] = 3 • 4 = 12

OpenStudy (anonymous):

this is the answer [(3 • 4) – 2] + (5 + 7)

OpenStudy (mathstudent55):

This is what I meant by understanding PEMDAS properly. Do operations of each line before the next line below. P - parentheses E - exponents MD - multiplications and divisions in the order they appear from left to right AS - additions and subtractions in the order they appear from left to right

OpenStudy (mathstudent55):

Many people think that PEMDAS means doing the operations in the order below, but that is incorrect. P - parentheses \ E - exponents \ M - multiplications \ D - divisions / The PEMDAS shown to the left is incorrect. A - additions / S -subtractions / The correct order is what I wrote in the previous response above with MD in one line and AS in one line.

OpenStudy (mathstudent55):

Yes, you are correct.